OpenStudy (anonymous):

rcos((theta)-pi/6)=2???? convert to rectangular?

OpenStudy (anonymous):

multiply the complex no z = x+iy by e^(-ipi/6)

OpenStudy (anonymous):

\[rcos (\Theta-\pi/6)=2\]

OpenStudy (amistre64):

cos(t-pi/6)=2

OpenStudy (anonymous):

you can always remember cos(a+b) = cosa * cosb - sina * sinb

OpenStudy (anonymous):

let the new complex no be z1 the x part of z1 is rcos(theta - pi/6)

OpenStudy (amistre64):

its just a straight line right?

OpenStudy (anonymous):

no clue:(

OpenStudy (anonymous):

@franc..wat happnd 2 the r in ur prev step?

OpenStudy (amistre64):

since cos only goes from -1 to 1; i think its a false statement

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

you can expand the cosine of sum of angles...then you would had something like ax+by = 2

OpenStudy (anonymous):

r is sqrt(x^2 + y^2) cos (t - pi/6) is ((root 3)x + y) / 2

OpenStudy (amistre64):

cos(doesnt matter) never = 2

OpenStudy (anonymous):

.did u get wht i wrote?

OpenStudy (anonymous):

remember r could be 2, then let theta = pi/6 ...done 2 = 2

OpenStudy (anonymous):

soo..in rectangular form..tht is....?

OpenStudy (anonymous):

whom r u asking?

OpenStudy (anonymous):

doesn't matter, the fact is you need to know what you have to do....have any idea?

OpenStudy (anonymous):

i told u guys

OpenStudy (anonymous):

this is not a site where you get answers, we discuss possible ways to find out the solution

OpenStudy (anonymous):

just rotate a complex no by pi/6 clockwise to get cos(theta - pi/6)

OpenStudy (anonymous):

ohhok..\[r=2/((\sqrt{3}/2)\cos \Theta-1/2(\sin \Theta))\]

OpenStudy (anonymous):

well, you got something (: now remember that r cos(theta) = x and r sin (theta) = y

OpenStudy (anonymous):

there is no more thinking in this problem

OpenStudy (anonymous):

z = r(cost + isint) let x = rcost y = rsint now ze^(-ipi/6) = r[cos(t-pi/6) + isin(t - pi/6)]

OpenStudy (anonymous):

if z= x+ iy \[ze^{-i \pi/6} = [\sqrt{3}x + y]/2 + i [\sqrt{3}y - x]/2\]

OpenStudy (anonymous):

getting it?

OpenStudy (anonymous):

somewhat...is der a way to do it without the i's?

OpenStudy (anonymous):

none tht i know

OpenStudy (anonymous):

so there u go if u compare the two eqns u find \[rcos(\theta- \pi /6) = [\sqrt{3}x + y]/2\]

OpenStudy (anonymous):

thats equal to 2

OpenStudy (anonymous):

ohhok..yeah i think i got it! thanks a lott

OpenStudy (anonymous):

√3x + y = 4

OpenStudy (anonymous):

no problem...just be comfy with complex nos and ull be fine